The goals of this proposal was to unify the entire process followed by RUL, ensure the security of non-print material, and allow our users to work independently. These procedures will be followed at most RUL sites.
There are two categories: Print and Non-print formats with accompanying materials, and Non-book formats that do not have accompanying materials.
A document was distributed to the group for review that highlighted some of the areas that will impact CM staff. Changes to an item record in the vol/copy folder, how accompanying materials will display in marc holdings statement, changes to a serials control record, and new charge and discharge procedures for accompanying materials were examined. The Group was asked to forward any questions or concerns to Andy.
Sandy from Camden commented that Camden has been doing this for a while with success and said that accompanying materials is a very small percentage of their collection. Alex from Dana also commented that the implementation of this at his site is proceeding smoothly.
The official name of the group has been changed to the Collection Management Group. "Functional" has been dropped from all of the group's names. We are now a subgroup of the Access Services Committee and report to that group through the Coordinator.
The draft of the purpose statement revision was accepted, with some revisions. The following revisions to the draft where made: CMG is now a sub-group of the Access Services Group, and the group is comprised of library staff from all three campuses. The purpose statement is a general statement depicting the work that what we do as a group. It is to be used as a guide for our current work and all future projects. It will be posted to the Access Services web page shortly.
Andy: Andy announced that Bill Lynch from Douglass would be joining the group in April/2006. You can contact Bill with CM questions at firstname.lastname@example.org. With regard to the Relative Humidity data loggers, as soon as Ian has a location for the information it would be a good idea for each unit to check periodically for fluctuation in RH and temperature in their library.
Krista/Kilmer: The sorting shelves for the library were recently moved to their new location at the Circulation desk on the sliding shelves. The microfilm collection will be shifted because the cabinets are running out of available space in some sections.
Paul/Alex & LSM: Paul is continuing to review shifting options with the GovDoc reference librarians at Alex, and is still shelving U.S. topographical maps at LSM.
Rob/Alexander: While working on 2A Rob noticed a smell and upon investigation found approximately 42 books that where moldy. At last, count 72 items had been found that where moldy. Searching continues at Alexander, along with investigation into the cause of this mold outbreak. He checked the humidity/temperature readings from Alexander and has found large fluctuations in the past months. Right now, searching has been contained on 2A, but it will be expanded to other floors in the future. He mentioned that the HVAC system is 40 years old.
Eddie/LSM: Eddie and CM staff continue to assist with the LSM Weeding Project, the goal is 9,000 remaining volumes to reach the weeding goal. Eddie continues with mini-shift of 3rd floor Bound Periodicals.
Mary/Alexander: Mary is currently working on searching for mold in the collection. They are instructing shelvers to keep their eyes open for anything that looks strange. Many students are taking the mold booklet home with them.
Shirley/Alexander: There has been an increase in the number of financial holds that patrons want to clear up.
Melanie/Math: The Math Library may finally be able to cancel the print version of the Lecture Notes in Computer Science. Melanie has completed the Active Titles portion of the Serials Inventory. She is currently working on inventory of Reference materials for a future transfer project.
Melody/CS-LSM: CS staff are making good progress on the LSM weeding project. Work will be starting soon on the Closed Titles portion of the Serials Inventory Project, after an acceptable start date is decided upon and training takes place.
Dean/Annex: During Rob's information sharing Dean stated that he has located books at the Annex that may be moldy. He will check with Farideh and Ian Bogus. He is working with Judy Gardner to calculate the fill rate of the annex's shelves per an ARL stacks spec kit. There was a water leak in the staff area so the humidifier has been turned off. Chad figured out how to calculate available shelving space using a spreed sheet.
Alex/Dana: Shifting lower level bound periodicals and conducting a minor shift of circulating books.
Andy asked the group to read the procedure he sent us to see if it is ok or should be revised. The group consensus is that bullet's work better than paragraphs for highlighting needed information. Andy said he would work on it and get it back to the group by the next meeting. Mary asked about what would we do if a book were wet? Andy said he would put a link in the web site to on how to handle wet books.
Andy is formatting the reports for each site and removing problem items. There are 712 items shadowed for all of NBL on the reports that will be searched (plus additional items for Camden and Dana). The reports will be distributed to CM staff, will be searched and all items found in the collections will be corrected. Items not found will be referred for investigation.
The Pest Awareness Sub-Group will be meeting on March 27th. Ian Bogus and Tim Corliss have been added to the sub-group. Rob has been putting together some information about pests and placed a document on the Douglass T drive so that the sub-group can look at it. Andy will be sending out e-mail to invite more group members to join the Best Practices Sub-group.
This discussion has been tabled to the next meeting.
Date: Wednesday, April 12, 2006
Time: 2:30 pm-4: 30 pm
Primary Location: Media Center Classroom, Kilmer Library, and Livingston
Campus Satellite Location(s): Special Collections Room: DANA # 454 (attendance: 1)
Conference Room: ROBESON # 290 (attendance: 1)